A body moving with uniform retardation covers covers 3km before it's speed is reduced half of value. What is another distance travelled before comes rest

Grade:9

1 Answers

Arun
5 years ago
 
let us suppose that its initial velocity = u
using equation of motion
v2=u2+2as
 
    (u/2)2 = u2 =2*a*3000 (value of final velocity is haif he initial taht is u/2 and s = 3 km =3000m)
\therefore (u2/4) – u=2*a*3000
\therefore (-3u2/4)=2*a*3000
\therefore – u2=(2*a*3000*4)/3
\therefore u= – 8000*a                      ….............................1
by appling the same equation when it comes to rest we get 
    02=(u/2)2 +2*a*s
\therefore – u2/4 =2*a*s
\therefore – u2/(4*2*a) = s
\therefore – u2/8a = s
after substituting values form 1 we get
\therefore – (-8000*a)/8a = s
\therefore s = 1000m = 1km